OpenStudy (binary3i):

what is the condition for any function to coinside with its inverse function other than at line y=x?

OpenStudy (anonymous):

it can be any identity function i spose

OpenStudy (anonymous):

?

OpenStudy (binary3i):

no

OpenStudy (anonymous):

sentence does not parse

OpenStudy (mathmate):

y=sqrt(x^2)

OpenStudy (anonymous):

The function must be one-to-one to have an inverse function (or it can be many-to-one with constraints)

OpenStudy (across):

Assume you have a function \(f(x)\) whose inverse is \(f^{-1}(x)\). If you define a function \(g(x)=f(f^{-1}(x))\), you'll end up with a function \(g(x)=x\) which lies on the line \(y=x\).\[\]

OpenStudy (mathmate):

Oops, y=x^2 doesn't quite work, because it resembles the absolute function, and the inverse does not coincide with the function.

OpenStudy (anonymous):

can you rephrase the question? you are getting several answers indicating that the question is not clear at all

OpenStudy (mathmate):

If f(f-1(x))=x, and f(x)=f-1(x), then f(f(x))=x would be a requirement.

OpenStudy (anonymous):

given a one to one function f, then \[f(a)=f^{-1}(b)\] if \[a=b\]

OpenStudy (binary3i):

i asked is their any point for f(x) which is also on \[ f^{-1}\] but not on line y=x if yes then what can you say about the function?